OpenStudy (jenny345g):

PLEASE HELP In a particular region of a national park, there are currently 330 deer, and the population is increasing at an annual rate of 11%. a. Write an exponential function to model the deer population. b. Explain what each value in the model represents. c. Predict the number of deer that will be in the region after five years.

OpenStudy (sunnnystrong):

Okay so recall that: \[A(n)=A _{0}(b)^n\] A(n)=Amount Ao=Initial B=Growth factor (1+percent of growth) n=Variable

OpenStudy (jenny345g):

Okay

OpenStudy (sunnnystrong):

@Jenny345G ... What do you think the exponential equation is? If Ao=330 & The percent of growth is 11% (.11)?

OpenStudy (jenny345g):

Uhm ) 330(1.11)^n ?

OpenStudy (sunnnystrong):

Yep!! :D Good job!

OpenStudy (sunnnystrong):

Now, what do these variables mean.... If \[f(x)=330(1.11)^x\]

OpenStudy (jenny345g):

The number of deer after x years?

OpenStudy (sunnnystrong):

Yep (: Last part is just asking you to compute f(5)

OpenStudy (jenny345g):

330*1.11^5 = 556 deer I think

OpenStudy (sunnnystrong):

Good job! (:

OpenStudy (jenny345g):

Thanks sunny!

OpenStudy (sunnnystrong):

Anytime :D Always happy to help!